what are three consecutive odd integers whose sum is equal to five more than two times the second largest number among the three numbers.
Answer by algebrahouse.com(1659)

x = first
x + 2 = second
x + 4 = third {odd integers increase by 2 each time}

x + x + 2 + x + 4 = 2(x + 2) + 5 {sum is five more than two times second largest}
3x + 6 = 2x + 4 + 5 {used distributive property and combined like terms}
3x + 6 = 2x + 9 {combined like terms on right side}
x = 3 {subtracted 2x and subtracted 6 from each side}
x + 2 = 5 {substituted 5, in for x, into x + 2}
x + 4 = 7 {substituted 3, in for x, into x + 4}

3,5, and 7 are the three consecutive odd integers

For more help from me, visit: www.algebrahouse.com